OpenStudy (anonymous):

Solve V = π r2h for h. Show all work. Help?

OpenStudy (anonymous):

welcome to openstudy

OpenStudy (anonymous):

r = 1.35 h = 12.7

OpenStudy (dan815):

hi

OpenStudy (dan815):

why dont you substitute those numbers into the formula and solve it

OpenStudy (tkhunny):

Divide by \(\pi r^{2}\). \(\left(\pi r^{2}\right)\) is attached to 'h' via multiplication. Detach it via division.

OpenStudy (anonymous):

@dan815 i clearly have done so already, but not coming up with the 'correct' answer according to the rest of the problem i'm inserting it into, hence why I'm seeing if im doing it wrong.

OpenStudy (anonymous):

@tkhunny hi ther, thanks, but i'm not trying to solvefor h, i'm subsituting the numbers i posted above. . . seem to be coming up incorrect according to the book/problem set

OpenStudy (dan815):

what is the object you are trying to find the integral of?

OpenStudy (dan815):

volume of* sorry

OpenStudy (dan815):

is it a cylinder?

OpenStudy (tkhunny):

?? "Solve V = π r2h for h" Perhaps we're speaking a different language? One might ask, why are nyou not trying to solve for h? Are you studying Significant Figures"? What answer are you giving that is being rejected?

OpenStudy (bradely):

step by step answers posted here http://www.mathskey.com/question2answer/2063/solve-v-%CF%80-r-2-h-for-h ask your Algebra1 homework questions at http://www.mathskey.com/question2answer/algebra-1 and get free math help. all the best

OpenStudy (anonymous):

its a beaker filled with a certain amount of liquid, all a part of a stream of different question s built off this one

OpenStudy (anonymous):

@tkhunny because i HAVE h already, as posted above, im trying to find V

OpenStudy (anonymous):

this is the complete problem : 5. Calculate the volume of the 50mL graduated cylinder using your measurements of diameter and height, using the formula V = ∏ r2h (r=½ diameter). This is your experimental value. Assuming the accepted value of the volume of the graduated cylinder is 50.00 mL, calculate the percent error of your volume calculation, using the following formula: Percent Error = /accepted value - experimental value/x 100 accepted value

OpenStudy (dan815):

lets compare answers i get for r = 1.35 h = 12.7 pi*2*r*h = 107.725212092

OpenStudy (tkhunny):

That's okay, but I'm really curious how this is consistent with "Solve V = π r2h for h". Anyway, it appears we ARE studying Significant Figures. Now answer the question. What value are you using that is being rejected?

OpenStudy (dan815):

that is area tho i think you mean pi * r^2 *h = 72.7145181618

OpenStudy (anonymous):

i'm getting 72.7145. . .but when i plug it into the next step in my problem it doesnt make any sense! im getting a negative percentage error

OpenStudy (dan815):

it says its pir2h so u shud use the other one 107.725212092

OpenStudy (dan815):

your question is wrong i think it shud say pir^2*2 not pi*r2h

OpenStudy (tkhunny):

Well, a negative value isn't possible if you are using the absolute values. Are you using the absolute values? Also, 72.71415 is no good. Use 72.7

OpenStudy (dan815):

pi*r^2 *h